### v1

@@ Abstract
Memory-based approaches have shown strong performance in Continual Relation Extraction (CRE). However, storing examples from previous tasks increases memory usage and raises privacy concerns. Recently, prompt-based methods have emerged as a promising alternative, as they do not rely on storing past samples. Despite this progress, current prompt-based techniques face several core challenges in CRE, particularly in accurately identifying task identities and mitigating catastrophic forgetting. Existing prompt selection strategies often suffer from inaccuracies, lack robust mechanisms to prevent forgetting in shared parameters, and struggle to handle both cross-task and within-task variations. In this paper, we propose WAVE++, a novel approach inspired by the connection between prefix-tuning and mixture of experts. Specifically, we introduce task-specific prompt pools that enhance flexibility and adaptability across diverse tasks while avoiding boundary-spanning risks; this design more effectively captures variations within each task and across tasks. To further refine relation classification, we incorporate label descriptions that provide richer, more global context, enabling the model to better distinguish among different relations. We also propose a training-free mechanism to improve task prediction during inference. Moreover, we integrate a generative model to consolidate prior knowledge within the shared parameters, thereby removing the need for explicit data storage. Extensive experiments demonstrate that WAVE++ outperforms state-of-the-art prompt-based and rehearsal-based methods, offering a more robust solution for continual relation extraction. Our code is publicly available at https://github.com/PiDinosauR2804/WAVE-CRE-PLUS-PLUS.
@@
Continual Relation Extraction , Rehearsal-Free Learning , Prompting Techniques , Mixture of Experts , Label Descriptions , Ensemble Learning
@@ 1 Introduction
Continual Relation Extraction (CRE) involves training models to progressively extract relationships between entity pairs across a sequence of tasks (Wang et al., 2019; Zhao et al., 2022; Nguyen et al., 2023; Le et al., 2025a). As a specialized domain of continual learning, CRE’s primary goal is to mitigate catastrophic forgetting (McCloskey and Cohen, 1989; Nguyen et al., 2019), where model performance deteriorates as the number of tasks increases. To address this issue, most CRE methods employ a memory buffer, which stores samples for revisiting previous tasks and preserving previously learned knowledge (Han et al., 2020; Zhao et al., 2022). Approaches utilizing this strategy are commonly referred to as rehearsal-based methods.
@@
While rehearsal-based methods have demonstrated notable success in mitigating catastrophic forgetting, they still exhibit significant limitations that call for more robust solutions. First, despite utilizing memory buffers, the representations of learned relations often degrade soon after training transitions to subsequent tasks (Caccia et al., 2021). Second, such methods violate a core principle of continual learning, as they necessitate ongoing access to datasets from earlier tasks. This reliance raises substantial concerns regarding data privacy and the considerable storage demands associated with large-scale memory buffers. Therefore, there is a pressing need to develop alternative approaches that minimize memory usage, thereby preserving the essence of continual learning while addressing privacy and scalability issues (Ke and Liu, 2022).
@@
Recent advances in continual learning take inspiration from prompt-based techniques in natural language processing, leading to a new class of methods that leverage learnable parameters, known as prompts, to guide pre-trained models on downstream tasks without requiring access to past data (Wang et al., 2022b, a, 2024; Tran et al., 2024a). In contrast to traditional memory-replay methods, these approaches do not rely on storing samples from previous tasks; instead, they insert small sets of auxiliary parameters to steer the training process. These prompts are adaptable to specific tasks, enabling continual relation extraction without the necessity of data replay.
@@
Despite these advantages, current prompt-based frameworks still exhibit several limitations. First, they do not fully alleviate catastrophic forgetting because many approaches rely on shared prompts, such as the Prompt Pool (Wang et al., 2022b), the General Prompt (G-Prompt) (Wang et al., 2022a), or a shared MLP classifier, thereby constraining their adaptability to varying task distributions. Second, the task-prediction mechanisms employed by these methods (Wang et al., 2022a, 2024) are prone to misclassification errors. Any mismatch between the prompts used during training and those employed at inference can degrade performance. Finally, prompt-based methods often fail to address both cross-task and within-task variability. For example, Wang et al. (2022b) introduce a common prompt pool, which can lead to multiple tasks sharing identical prompts and diminish cross-task diversity. This limitation is especially prominent in continual relation extraction, where instances from different relation classes may present almost identical contexts, illustrated by examples such as:
@@
In alignment with these approaches, Le et al. (2024a) investigate the connection between Prefix-tuning (Li and Liang, 2021), a widely adopted prompt-based technique, and Mixture of Experts (MoE) models (Jacobs et al., 1991; Jordan and Jacobs, 1994). Their findings reveal that self-attention can be interpreted as embedding multiple MoE models, and that implementing prefix-tuning is akin to appending new prefix experts to these pre-trained MoE architectures, thereby enabling the fine-tuning of underlying representations.
Building on this insight, we propose WAVE++ (Within-Task Variance Awareness for CRE), a prompt-based method designed to address the limitations outlined above. Rather than employing a single prompt pool across all tasks, WAVE++ assigns a dedicated Prompt Pool for each task, thereby improving plasticity in response to task distribution shifts and better capturing task-specific characteristics. Additionally, WAVE++ integrates label descriptions of relations to learn global, relation-specific contexts, ensuring that essential characteristics of each relation are preserved.
To mitigate catastrophic forgetting within the shared parameters, WAVE++ incorporates generative models that produce instructed latent data representations for replay. This strategy guards against representation drift without requiring the storage of extensive raw data, offering a more efficient and comprehensive alternative to conventional replay-based methods. Further, although recent approaches often rely on a classifier head for task prediction (Wang et al., 2024; Le et al., 2025a), WAVE++ removes this dependency by selecting the relevant prompt pool through a simple voting mechanism. Experimental results demonstrate that WAVE++ achieves state-of-the-art performance, outperforming existing prompt-based and rehearsal-based baselines.
@@
    Through extensive experiments, we show that WAVE++ outperforms state-of-the-art prompt-based and rehearsal-based methods, validating both its effectiveness and versatility.
Figure 1: Overall framework of WAVE++. To mitigate forgetting across tasks, each $k$-th task is associated with its own prompt pool $\mathcal{P}_{k}$, rather than relying on a single, shared prompt pool as in L2P. In addition, we employ representation generators to synthesize information from previously learned tasks, thereby reinforcing the relation classifier’s capacity to retain accumulated knowledge.
@@ 2 Background and Related Work
In this section, we briefly describe the relevant background and discuss existing literature on continual relation extraction in Section 2.1. We then examine recent advancements in prompt-based approaches, followed by an overview of mixture of experts models in Section 2.2 and Section 2.3, respectively.
@@ 2.1 Continual Relation Extraction
Figure 2: Comparison with rehearsal-based methods. Unlike approaches that rely on a rehearsal buffer, L2P employs a single backbone model and a Prompt Pool to store task-specific knowledge, thereby eliminating the need for explicit rehearsal to prevent forgetting. L2P further adapts to each instance by selecting and updating prompts from the pool on a per-instance basis.
Relation extraction is a subfield of continual learning Hai et al. (2024); Phan et al. (2022); Van et al. (2022) and continual information extraction Le et al. (2024b); Dao et al. (2024); Tran et al. (2024b). It is a fundamental task in natural language processing (NLP) that aims at identifying semantic relationships between entities in unstructured text. By capturing how entities relate to one another, relation extraction facilitates deeper understanding of textual data and serves as a foundation for numerous downstream applications. Continual relation extraction (CRE) is a special branch of relation extraction that addresses the challenge of incrementally learning new relational patterns from incoming tasks while retaining the ability to recognize relations learned in previous tasks.
@@
Formally, CRE involves learning sequentially from a series of tasks $\{\mathcal{T}_{1},...,\mathcal{T}_{k}\}$, where $\mathcal{T}_{t}$ denotes the $t$-th task. Similar to traditional supervised classification frameworks (Ji et al., 2020), each task $\mathcal{T}_{t}$ is associated with a dataset $\mathcal{D}_{t}=\{({\bm{x}}^{t}_{i},y^{t}_{i})\ |\ y^{t}_{i}\in\mathcal{R}_{t}\}_{i=1}^{\mathcal{N}_{t}}$ of $\mathcal{N}_{t}$ input–label pairs and a distinct set of relations $\mathcal{R}_{t}$. Notably, these relation sets are mutually exclusive across tasks, such that $\mathcal{R}_{i}\bigcap\mathcal{R}_{j}=\emptyset$ for all $i\neq j$ and $i,j\in\{1,\dots,k\}$. Furthermore, once the model completes training on a given task $t$, it can no longer access the data from that task when learning subsequent tasks.
As training proceeds, the model must maintain its performance on all relations encountered so far, represented as ^$\widehat{\mathcal{R}}_{t}=\bigcup_{i=1}^{t}\mathcal{R}_{i}$. A primary challenge under these constraints is catastrophic forgetting, wherein performance on earlier tasks deteriorates as the model adapts to new tasks. (McCloskey and Cohen, 1989; Nguyen et al., 2019).
@@
Many strategies have been explored to address this challenge. For instance, EA-EMR (Wang et al., 2019) introduces a technique that combines memory replay and embedding alignment to alleviate catastrophic forgetting. RP-CRE (Cui et al., 2021) employs a memory network to refine sample embeddings with relation prototypes. Similarly, CRE-DAS (Zhao et al., 2023) introduces memory-insensitive relation prototypes and memory augmentation to mitigate overfitting. CDec+ACA (Xia et al., 2023) proposes a classifier decomposition framework to address representation biases by promoting robust representation learning while preserving previously acquired knowledge. CREST (Le et al., 2024c) approaches the problem by training multiple objectives through sequential multi-task learning. However, most existing approaches for CRE rely on the use of a memory buffer to store samples of previously encountered relations, which raises critical concerns regarding both memory overhead and privacy (see Figure 2, left).
@@
where the first term refers to the model’s ability to determine the appropriate task, and the second term represents the model’s ability to accurately predict the label within that task. It has been shown that improving these two components is both a necessary and sufficient condition to enhance overall continual learning performance. This framework provides a solid foundation for designing more effective rehearsal-free continual learning algorithms (Wang et al., 2024; Le et al., 2024a). In this work, we build on this theoretical framework by proposing methods to improve both TII and WTP (see Section 4 for details).
@@ 2.2 Prompt-based Approaches
Owing to robust generalization capabilities, recent studies have increasingly leveraged pre-trained models to mitigate catastrophic forgetting, demonstrating impressive performance gains (Wang et al., 2022b, 2023a, 2023b; Zhou et al., 2024). In this work, we adopt a pre-trained BERT model (Devlin, 2018) as our base architecture, in line with previous research (Xia et al., 2023; Zhao et al., 2023). However, conventional fine-tuning of these models, which typically involves updating all parameters, poses substantial computational challenges.
To address these , Parameter Efficient Fine-tuning (PEFT) strategies have been proposed as an alternative paradigm (Houlsby et al., 2019; Hu et al., 2021; Lester et al., 2021; Li and Liang, 2021; Le et al., 2025b). PEFT focuses on designing fine-tuning methods that update only a small subset of parameters while leaving the rest of the model unchanged. By doing so, it significantly reduces the computational overhead associated with full fine-tuning. Among these techniques, prompt-tuning (Lester et al., 2021) and prefix-tuning (Li and Liang, 2021) have emerged as simple yet effective methods. Both approaches append learnable prompt tokens to the input—either at the input layer alone or also within intermediate representations—to serve as instructions for the pre-trained Transformer model to adapt to downstream tasks. Specifically, the Transformer architecture comprises multiple consecutive multi-head self-attention (MSA) layers (Vaswani, 2017), which are defined as follows:
@@
Building on the attention mechanism, prompt-tuning and prefix-tuning introduce learnable prompt parameters at the inputs of these MSA layers, enabling efficient fine-tuning of the pre-trained Transformer model.
Definition 2.2 (Prompt-tuning).
Let ${\mathbf{P}}=\left[{\bm{p}}_{1},\dots,{\bm{p}}_{L}\right]^{\top}\in\mathbb{R}^{L\times d}$ be the prompt parameters, where $L$ denotes the prompt length. Prompt-tuning appends the same prompt parameters to the query, key, and value matrices:
$f^{\mathrm{Pro-T}}_{\mathrm{prompt}}({\mathbf{P}},{\mathbf{X}}^{Q},{\mathbf{X}}^{K},{\mathbf{X}}^{V})$ $=\mathrm{MSA}\left(\begin{bmatrix}{\mathbf{P}}\\ {\mathbf{X}}^{Q}\end{bmatrix},\begin{bmatrix}{\mathbf{P}}\\ {\mathbf{X}}^{K}\end{bmatrix},\begin{bmatrix}{\mathbf{P}}\\ {\mathbf{X}}^{V}\end{bmatrix}\right)\in\mathbb{R}^{(N+L)\times d}.$ (3)
@@
Unlike prompt-tuning, prefix-tuning does not prepend prompts to the query matrix, thereby preserving the original input sequence dimension. In this study, we focus on prefix-tuning for prompt implementation. Note that only the prompt parameters ${\mathbf{P}}$ are trained, while the parameters of the pre-trained Transformer backbone including $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$, and $W^{O}$ remain frozen throughout training.
Inspired by these prompting techniques, several recent methods have successfully incorporated them into continual learning, providing a flexible and robust strategy to mitigate catastrophic forgetting without the need to store historical data, such as L2P (Wang et al., 2022b), DualPrompt (Wang et al., 2022a), and HiDe-Prompt (Wang et al., 2024). A more detailed discussion of these methods is presented in Section 3.

### v2

@@ Abstract
Memory-based approaches have shown strong performance in Continual Relation Extraction (CRE). However, storing examples from previous tasks increases memory usage and raises privacy concerns. Recently, prompt-based methods have emerged as a promising alternative, as they do not rely on storing past samples. Despite this progress, current prompt-based techniques face several core challenges in CRE, particularly in accurately identifying task identities and mitigating catastrophic forgetting. Existing prompt selection strategies often suffer from inaccuracies, lack robust mechanisms to prevent forgetting in shared parameters, and struggle to handle both cross-task and within-task variations. In this paper, we propose WAVE++, a novel approach inspired by the connection between prefix-tuning and mixture of experts. Specifically, we introduce task-specific prompt pools that enhance flexibility and adaptability across diverse tasks while avoiding boundary-spanning risks; this design more effectively captures both within-task and cross-task variations. To further refine relation classification, we incorporate label descriptions that provide richer, more global context, enabling the model to better distinguish among different relations. We also propose a training-free mechanism to improve task prediction during inference. Moreover, we integrate a generative model to consolidate prior knowledge within the shared parameters, thereby removing the need for explicit data storage. Extensive experiments demonstrate that WAVE++ outperforms state-of-the-art prompt-based and rehearsal-based methods, offering a more robust solution for continual relation extraction. Our code is publicly available at https://github.com/PiDinosauR2804/WAVE-CRE-PLUS-PLUS.
@@
Continual Relation Extraction , Prompting Techniques , Mixture of Experts , Label Descriptions , Ensemble Learning
@@ 1 Introduction
Continual Relation Extraction (CRE) involves training models to progressively extract relationships between entity pairs across a sequence of tasks (Wang et al., 2019; Zhao et al., 2022; Nguyen et al., 2023; Le et al., 2025a). As a specialized domain of continual learning, CRE’s primary goal is to mitigate catastrophic forgetting (McCloskey and Cohen, 1989; Nguyen et al., 2019), where model performance deteriorates as the number of tasks increases. To address this issue, most CRE methods employ a memory buffer, which stores samples to revisit previous tasks and preserve learned knowledge (Han et al., 2020; Zhao et al., 2022). Approaches utilizing this strategy are commonly referred to as rehearsal-based methods.
@@
Although rehearsal-based methods have shown notable success in mitigating catastrophic forgetting, they still exhibit significant limitations that demand more robust solutions. First, despite the use of memory buffers, the representations of learned relations often degrade shortly after training shifts to subsequent tasks (Caccia et al., 2021). Second, these methods violate a core principle of continual learning, by necessitating ongoing access to data from previous tasks. This reliance raises substantial concerns about data privacy and the considerable storage demands of large-scale memory buffers. A pressing need therefore exists for alternative approaches that minimize memory usage, thereby preserving the principles of continual learning while addressing critical privacy and scalability challenges (Ke and Liu, 2022).
@@
Recent advances in continual learning take inspiration from prompt-based techniques in natural language processing, leading to a new class of methods. These methods leverage learnable parameters, known as prompts, to guide pre-trained models on downstream tasks without requiring access to past data (Wang et al., 2022b, a, 2024). Unlike traditional memory-replay methods, these approaches do not rely on storing samples from previous tasks; instead, they insert small sets of auxiliary parameters to steer the training process. These prompts are adaptable to specific tasks, enabling continual relation extraction without the need for data replay.
@@
Despite these advantages, current prompt-based frameworks still exhibit several limitations. First, they do not fully alleviate catastrophic forgetting because many approaches rely on shared components, such as the Prompt Pool (Wang et al., 2022b), the General Prompt (G-Prompt) (Wang et al., 2022a), or a shared MLP classifier, thereby constraining their adaptability to varying task distributions. Second, the task-prediction mechanisms employed by these methods (Wang et al., 2022a, 2024) are prone to misclassification errors, as any mismatch between the prompts used during training and those employed at inference can degrade performance. Finally, prompt-based methods often fail to address both cross-task and within-task variability. For example, Wang et al. (2022b) introduce a single, shared prompt pool, which can result in inputs from different relations using the same or very similar prompts. This limitation is especially prominent in continual relation extraction, where instances from different relation classes may present almost identical contexts, as illustrated by the following examples:
@@
When such instances share the same prompts, their final hidden representations may collapse—becoming too similar for the relation classifier head to distinguish between them effectively.
In alignment with these approaches, Le et al. (2024a) investigate the connection between Prefix-tuning (Li and Liang, 2021), a widely adopted prompt-based technique, and Mixture of Experts (MoE) models (Jacobs et al., 1991; Jordan and Jacobs, 1994). Their findings reveal that self-attention can be interpreted as containing multiple MoE models, and that implementing prefix-tuning is equivalent to adding new prefix experts to these pre-trained MoE architectures, thereby enabling the fine-tuning of underlying representations.
Building on this insight, we propose WAVE++ (Within-Task Variance Awareness for CRE), a prompt-based method designed to address the prior limitations. Rather than employing a single prompt pool across all tasks, WAVE++ assigns a dedicated prompt pool for each task, thereby improving plasticity in response to task distribution shifts and capturing task-specific characteristics. Additionally, WAVE++ integrates label descriptions of relations to learn global, relation-specific contexts, ensuring that essential characteristics of each relation are preserved.
To mitigate catastrophic forgetting in the shared parameters, WAVE++ incorporates generative models that produce instructed latent data representations for replay. This strategy guards against representation drift without requiring the storage of extensive raw data, offering a more efficient and comprehensive alternative to conventional replay-based methods. Furthermore, although recent approaches often rely on a classifier head for task prediction (Wang et al., 2024; Le et al., 2025a), WAVE++ removes this dependency by selecting the relevant prompt pool through a simple voting mechanism. Experimental results show that WAVE++ achieves state-of-the-art performance, outperforming existing prompt-based and rehearsal-based baselines.
@@
    We conduct extensive experiments demonstrating that WAVE++ outperforms state-of-the-art prompt-based and rehearsal-based methods, validating both its effectiveness and versatility.
Figure 1: Overall framework of WAVE++. To mitigate forgetting across tasks, each $k$-th task is associated with its own prompt pool $\mathcal{P}_{k}$, rather than relying on a single, shared prompt pool as in L2P. In addition, we employ representation generators to synthesize information from previously learned tasks, thereby reinforcing the capacity of the relation classifier to retain accumulated knowledge.
@@ 2 Background and Related Work
In this section, we briefly review the relevant background of continual relation extraction in Section 2.1. We then examine recent advancements in prompt-based approaches, followed by an overview of MoE models in Section 2.2 and Section 2.3, respectively.
@@ 2.1 Continual Relation Extraction
Figure 2: Comparison with rehearsal-based methods. Unlike approaches that rely on a rehearsal buffer, L2P employs a single backbone model and a prompt pool to store task-specific knowledge, thereby eliminating the need for explicit rehearsal to prevent catastrophic forgetting. L2P further adapts to each instance by selecting and updating prompts from the pool on a per-instance basis.
@@
Continual relation extraction (CRE) is a special subfield of relation extraction that addresses the challenge of incrementally learning new relational patterns from incoming tasks while preserving knowledge of previously learned relations. Formally, CRE involves learning sequentially from a series of tasks $\{\mathcal{T}_{1},...,\mathcal{T}_{k}\}$, where $\mathcal{T}_{t}$ denotes the $t$-th task. Similar to traditional supervised classification frameworks (Ji et al., 2020), each task $\mathcal{T}_{t}$ is associated with a dataset $\mathcal{D}_{t}=\{({\bm{x}}^{t}_{i},y^{t}_{i})\ |\ y^{t}_{i}\in\mathcal{R}_{t}\}_{i=1}^{\mathcal{N}_{t}}$ containing $\mathcal{N}_{t}$ input–label pairs and a distinct set of relations $\mathcal{R}_{t}$. Notably, these relation sets are mutually exclusive across tasks, such that $\mathcal{R}_{i}\bigcap\mathcal{R}_{j}=\emptyset$ for all $i,j\in\{1,\dots,k\}$ where $i\neq j$. Furthermore, once the model completes training on a given task $t$, it can no longer access the data from that task when learning subsequent tasks.
@@
As training proceeds, the model must maintain its performance on all relations encountered so far, represented as $\widehat{\mathcal{R}}_{t}=\bigcup_{i=1}^{t}\mathcal{R}_{i}$. A primary challenge under these constraints is catastrophic forgetting, wherein performance on earlier tasks deteriorates as the model adapts to new ones (McCloskey and Cohen, 1989; Nguyen et al., 2019). Many strategies have been explored to address this challenge. For instance, EA-EMR (Wang et al., 2019) combines memory replay and embedding alignment to alleviate catastrophic forgetting. RP-CRE (Cui et al., 2021) employs a memory network to refine sample embeddings with relation prototypes. Similarly, CRE-DAS (Zhao et al., 2023) introduces memory-insensitive relation prototypes and memory augmentation to mitigate overfitting. CDec+ACA (Xia et al., 2023) proposes a classifier decomposition framework to address representation biases by learning robust representations while preserving previously acquired knowledge. However, most existing approaches for CRE rely on a memory buffer to store samples of previously encountered relations, which raises critical concerns regarding memory overhead and privacy (see Figure 2, left).
@@
where the first term refers to the model’s ability to determine the appropriate task, and the second term represents its ability to accurately predict the label within that task. It has been shown that improving these two components is both a necessary and sufficient condition for enhancing overall continual learning performance. This framework provides a solid foundation for designing more effective rehearsal-free continual learning algorithms (Wang et al., 2024; Le et al., 2024a). In this work, we build upon this theoretical framework by proposing methods to improve both TII and WTP (see Section 4 for details).
@@ 2.2 Prompt-based Approaches
Owing to robust generalization capabilities, recent studies have increasingly leveraged pre-trained models to mitigate catastrophic forgetting, achieving impressive performance gains (Wang et al., 2022b, 2023a, 2023b; Zhou et al., 2024). Following previous research (Xia et al., 2023; Zhao et al., 2023), we adopt a pre-trained BERT model (Devlin, 2018) as our base architecture. However, the conventional approach of fine-tuning all model parameters presents substantial computational challenges. To address this challenge, Parameter Efficient Fine-Tuning (PEFT) strategies have been proposed as an alternative paradigm (Houlsby et al., 2019; Hu et al., 2021; Lester et al., 2021; Li and Liang, 2021; Le et al., 2025b). PEFT focuses on designing fine-tuning methods that update only a small subset of parameters while leaving the rest of the model unchanged. By doing so, it significantly reduces the computational overhead associated with full fine-tuning. Among these techniques, prompt-tuning (Lester et al., 2021) and prefix-tuning (Li and Liang, 2021) have emerged as simple yet effective methods. Both approaches append learnable prompt tokens to the input, either at the input layer alone or within intermediate representations, to serve as instructions for the pre-trained Transformer model to adapt to downstream tasks. In this work, we adopt prefix-tuning as our prompting approach. Specifically, the Transformer architecture comprises multiple consecutive multi-head self-attention (MSA) layers (Vaswani, 2017), which are defined as follows:
@@
Building on the attention mechanism, prefix-tuning introduces learnable prompt parameters at the inputs of these MSA layers, enabling efficient fine-tuning of the pre-trained Transformer model.
@@
By not prepending prompts to the query matrix, prefix-tuning preserves the original input sequence dimension. Importantly, only the prompt parameters $\mathbf{P}$} are trained, while the parameters of the pre-trained Transformer backbone including $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$, and $W^{O}$ remain frozen. Inspired by these prompting techniques, several continual learning methods, such as L2P (Wang et al., 2022b), DualPrompt (Wang et al., 2022a), and HiDe-Prompt (Wang et al., 2024), mitigate catastrophic forgetting without requiring historical data. These methods are discussed in greater detail in Section 3.